# Predictive Uncertainty Quantification via Risk Decompositions for Strictly Proper Scoring Rules

### v2

@@ Appendix A Derivations of different risks with proper scoring rules
Appendix B Derivation of risks for specific choices of scoring rules
@@
B.1 Log score (Cross-Entropy)
@@
B.2 Quadratic score (Brier Score)
@@ B.3 Zero-one score
$\ell(\eta,i)=\langle\mathbb{I}[k=\operatorname*{arg\,max}_{j}\eta_{j}]\,,\,\eta\rangle-\mathbb{I}[i=\operatorname*{arg\,max}_{j}\eta_{j}]-\max_{k}\eta_{k}+1=\\ \max_{k}\eta_{k}-\mathbb{I}[i=\operatorname*{arg\,max}_{j}\eta_{j}]-\max_{k}\eta_{k}+1=1-\mathbb{I}[i=\operatorname*{arg\,max}_{j}\eta_{j}],$
@@ B.5 Negative log score
Appendix C Approximations
In this section, we provide derivations for different types of Bayesian approximations. First, let us note, that Bayes risk depends only on the ground truth density (hence has only one argument), while Total and Excess risks are functions of both ground truth density and our approximation (they have two arguments).
Given a set of models (ensemble or samples from a posterior distribution over model parameters), we can use these approximation ideas for both of the arguments.
Total risk.
Let us start with Total risk. Since $\text{R}_{\text{Tot}}$ depends linearly on $\eta$, both approximations of the ground truth density lead to the same result:
$\tilde{\text{R}}_{\text{Tot}}(\theta\mid x)=\langle G^{\prime}(\hat{\eta}_{\theta})\,,\,\hat{\eta}_{\theta}\rangle-G(\hat{\eta}_{\theta})-\langle G^{\prime}(\hat{\eta}_{\theta})\,,\,\hat{\eta}_{D_{tr}}\rangle,$ (6)
where $\hat{\eta}_{D_{tr}}(x)=\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}\hat{\eta}_{\tilde{\theta}}(x)$ denotes posterior predictive distribution.
We also have a second argument, for which we also can incorporate this Bayesian reasoning. To do so, we will use an ensemble. Again, two approximation ideas, which lead to the following:
$\tilde{\text{R}}_{\text{Tot}}^{(1)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}\Bigl{(}\langle G^{\prime}(\hat{\eta}_{\theta})\,,\,\hat{\eta}_{\theta}\rangle-G(\hat{\eta}_{\theta})-\langle G^{\prime}(\hat{\eta}_{\theta})\,,\,\hat{\eta}_{D_{tr}}\rangle\Bigr{)},$
@@ Appendix C Approximations
$\tilde{\text{R}}_{\text{Tot}}^{(2)}(x)=\langle G^{\prime}(\hat{\eta}_{D_{tr}})\,,\,\hat{\eta}_{D_{tr}}\rangle-G(\hat{\eta}_{D_{tr}})-\langle G^{\prime}(\hat{\eta}_{D_{tr}})\,,\,\hat{\eta}_{D_{tr}}\rangle=-G(\hat{\eta}_{D_{tr}}),$
where by superscripts we denote the index of approximation idea.
As we will see below, the second case effectively corresponds to the situation with zero Excess risk (no epistemic uncertainty).
Bayes risk.
For Bayes risks, these two approximation ideas lead to the following results:
$\tilde{\text{R}}_{\text{Bayes}}^{(1)}(x)=-\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}G(\hat{\eta}_{\tilde{\theta}})\quad\text{and}\quad\tilde{\text{R}}_{\text{Bayes}}^{(2)}(x)=-G(\hat{\eta}_{D_{tr}}).$
However, it is not clear, which of these approximations is better. To investigate it, we assume that there exists a vector of true parameters $\theta^{*}$, for which $\eta(x)=p(y\mid x)=p(y\mid x,\theta^{*})$ for any input point $x$. For example for neural networks, which are flexible enough to represent any function, this is a mild assumption.
Note, that according to equation (4), Bayes risk is concave. Using $g(.)=-\mathbb{E}_{x}G(.)$, the following follows from Jensen’s inequality:
$\mathbb{E}_{p(\theta\mid D_{tr})}g(\hat{\eta}_{\theta})\leq g(\mathbb{E}_{p(\theta\mid D_{tr})}\hat{\eta}_{\theta}).$
$g(\eta)=g\bigl{(}p(y\mid x,\theta^{*})\bigr{)}\leq g(\hat{\eta}_{\theta}),$
and in particular:
$g(\eta)\leq\mathbb{E}_{p(\theta\mid D_{tr})}g(\hat{\eta}_{\theta}).$
Hence, we have the following:
$g(\eta)\leq\mathbb{E}_{p(\theta\mid D_{tr})}g(\hat{\eta}_{\theta})\leq g(\mathbb{E}_{p(\theta\mid D_{tr})}\hat{\eta}_{\theta}).$ (7)
From equation (7) we see that for estimating Bayes risk, it is beneficial to use the first approximation idea, as it leads to a tighter upper-bound on a true risk, under the assumption that true parameter $\theta^{*}$ belongs to the set of possible parameters of the model.
Excess risk.
For Excess risk these two approximation ideas for the ground truth distribution lead to the following results:
$\tilde{\text{R}}_{\text{Exc}}^{(1)}(\theta\mid x)=\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}~{}\|~{}\hat{\eta}_{\theta}\Bigr{)}\quad\text{and}\quad\tilde{\text{R}}_{\text{Exc}}^{(2)}(\theta\mid x)=D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}~{}\|~{}\hat{\eta}_{\theta}\Bigr{)}.$
Furthermore, using results of [35], we can rewrite the first approximation as follows:
$\tilde{\text{R}}_{\text{Exc}}^{(1)}(\theta\mid x)=D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}~{}\|~{}\hat{\eta}_{\theta}\Bigr{)}+\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}~{}\|~{}\hat{\eta}_{D_{tr}}\Bigr{)}=\tilde{\text{R}}_{\text{Exc}}^{(2)}(\theta\mid x)+\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}~{}\|~{}\hat{\eta}_{D_{tr}}\Bigr{)}.$
We see that the first approximation contains the second one, and as Bregman divergence is non-negative, $\tilde{\text{R}}_{\text{Exc}}^{(1)}(\theta\mid x)\geq\tilde{\text{R}}_{\text{Exc}}^{(2)}(\theta\mid x)$ (which also follows from the convexity of Bregman in its first argument). Moreover, the difference between estimates is equal to $\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}\|\hat{\eta}_{D_{tr}}\Bigr{)}$, which is known [2] as a Bregman Information. A specific case of Bregman information for Log score (see Table 1) is BALD [10, 16], a well-known epistemic uncertainty measure. We see, that BALD is only a part of the epistemic uncertainty estimate. Worth noting, that in recent work [46, 38] concerns regarding this expression as a proper measure of epistemic uncertainty were raised.
Furthermore, similarly to the Total risk estimation, we can use ensembles to incorporate the Bayesian approximation idea for the second argument. Thus, we end up with four different estimates of Excess risk:
  • Expected Pairwise Bregman Divergence (EPBD):
    $\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)=\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}\mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}~{}\|~{}\hat{\eta}_{\theta}\Bigr{)}.$
    Note, that since KL divergence is a special case of Bregman divergence, Expected Pairwise KL (EPKL [30, 38]) is one of the special cases of this Excess risk estimate.
  • Bregman Information (BI):
    $\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x)=\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}~{}\|~{}\hat{\eta}_{D_{tr}}\Bigr{)},$
    which special case is BALD [10, 16].
  • Reverse Bregman Information (RBI):
    $\tilde{\text{R}}_{\text{Exc}}^{(2,1)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}~{}\|~{}\hat{\eta}_{\theta}\Bigr{)}.$
    Its special case for Log score is known as Reverse Mutual Information [30].
  • Finally, we obtain
@@
However, it is not clear, what estimate of the Excess risk we should use. Indeed, neither of these estimates are upper nor lower bounds for the true Excess risk. This is because contrary to Bayes risk, we don’t have any idea if Excess risk with ground truth $\eta$ reaches any extreme. For another explanation, see Figure 3 and discussion in Section 7.
@@
Bottom. $\tilde{\text{R}}_{\text{Tot}}$ overestimates $\text{R}_{\text{Tot}}$, $\tilde{\text{R}}_{\text{Bayes}}^{(1)}$ better estimates $\text{R}_{\text{Bayes}}$, and $\tilde{\text{R}}_{\text{Exc}}^{(2)}$ better estimates $\text{R}_{\text{Exc}}$. We see, that for different estimates of $\text{R}_{\text{Tot}}$, we have different best approximations for $\text{R}_{\text{Exc}}$. See discussion in Section 7.
@@
Appendix D Relations between the estimates
@@
Let us start with $\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)$. Using results of [35], we have:
$\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}\|\hat{\eta}_{\theta}\Bigr{)}=\\ \mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}\|\hat{\eta}_{\theta}\Bigr{)}+\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{\tilde{\theta}}\|\hat{\eta}_{D_{tr}}\Bigr{)}=\tilde{\text{R}}_{\text{Exc}}^{(2,1)}(x)+\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x).$
@@
$\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)\geq\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x)\geq\tilde{\text{R}}_{\text{Exc}}^{(2,2)}(x)=0$

### v3

# From Risk to Uncertainty: Generating Predictive Uncertainty Measures via Bayesian Estimation

@@ Appendix C Derivations of different risks with proper scoring rules
Appendix D Derivation of risks for specific choices of scoring rules
@@
D.1 Log score (Cross-Entropy)
@@
D.2 Quadratic score (Brier Score)
@@ D.3 Zero-one score
$\ell(\eta,i)=\langle\mathbb{I}[k=\operatorname*{arg\,max}_{j}\eta_{j}]\,,\,\eta\rangle-\mathbb{I}[i=\operatorname*{arg\,max}_{j}\eta_{j}]-\max_{k}\eta_{k}+1=\\ \max_{k}\eta_{k}-\mathbb{I}[i=\operatorname*{arg\,max}_{j}\eta_{j}]-\max_{k}\eta_{k}+1=1-\mathbb{I}[i=\operatorname*{arg\,max}_{j}\eta_{j}]=\mathbb{I}[i\neq\operatorname*{arg\,max}_{j}\eta_{j}],$
@@ D.5 Negative log score
Appendix E Is there the best approximation?
We discussed the choice of the approximation for Bayes risk in the main part of the paper. Here, we discuss whether there is a best choice of Excess risk. If we know it, we can choose the best Total risk.
Recall that for Excess risk there are 9 possible options:
  • Expected Pairwise Bregman Divergence (EPBD):
    $\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)=\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}\mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\eta_{\tilde{\theta}}~{}\|~{}\eta_{\theta}\Bigr{)}.$
    Note, that since KL divergence is a special case of Bregman divergence, Expected Pairwise KL (EPKL (Malinin & Gales, 2021; Schweighofer et al., 2023a)) is one of the special cases of this Excess risk estimate.
  • Bregman Information (BI):
    $\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x)=\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\eta_{\tilde{\theta}}~{}\|~{}\hat{\eta}_{D_{tr}}\Bigr{)},$
    which special case is BALD (Gal et al., 2017; Houlsby et al., 2011).
  • Reverse Bregman Information (RBI):
    $\tilde{\text{R}}_{\text{Exc}}^{(2,1)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}~{}\|~{}\eta_{\theta}\Bigr{)}.$
    Its special case for Log score is known as Reverse Mutual Information (Malinin & Gales, 2021).
  • Modified Bregman Information (MBI):
    $\tilde{\text{R}}_{\text{Exc}}^{(1,3)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\eta_{\theta}~{}\|~{}\bar{\eta}\Bigr{)}.$
    It is similar to Bregman Information, but the deviation is computed from the “central prediction”, not central label (BMA).
  • Modified Reverse Bregman Information (MRBI):
    $\tilde{\text{R}}_{\text{Exc}}^{(3,1)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\bar{\eta}~{}\|~{}\eta_{\theta}\Bigr{)}.$
    This has similar structure to Reverse Bregman Information (RBI). Again, the deviation here is computed from another “central” prediction.
  • Forward Bias term:
    $\tilde{\text{R}}_{\text{Exc}}^{(2,3)}(x)=D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}~{}\|~{}\bar{\eta}\Bigr{)}.$
    It represents a Bregman divergence between two different Bayesian estimates and can be viewed as a bias.
  • Reverse Bias term:
    $\tilde{\text{R}}_{\text{Exc}}^{(3,2)}(x)=D_{G}\Bigl{(}\bar{\eta}~{}\|~{}\hat{\eta}_{D_{tr}}\Bigr{)}.$
    This applies similar reasoning to the Forward Bias Term.
  • Finally, we obtain two similar constant measures:
    $\tilde{\text{R}}_{\text{Exc}}^{(3,3)}(x)=D_{G}\Bigl{(}\bar{\eta}~{}\|~{}\bar{\eta}\Bigr{)}=0,$
@@ Appendix E Is there the best approximation?
However, it is not clear, what estimate of the Excess risk we should use. Indeed, neither of these estimates are upper nor lower bounds for the true Excess risk. This is because contrary to Bayes risk, we don’t have any idea if Excess risk with ground truth $\eta$ reaches any extreme. For another explanation, see Figure 2 and discussion in Appendix B. For simplicity, we will consider only approximations with strategies (1) and (2), as for them we know, at least, which estimate of Bayes risk is better.
@@
Bottom. $\tilde{\text{R}}_{\text{Tot}}$ overestimates $\text{R}_{\text{Tot}}$, $\tilde{\text{R}}_{\text{Bayes}}^{(1)}$ better estimates $\text{R}_{\text{Bayes}}$, and $\tilde{\text{R}}_{\text{Exc}}^{(2)}$ better estimates $\text{R}_{\text{Exc}}$. We see, that for different estimates of $\text{R}_{\text{Tot}}$, we have different best approximations for $\text{R}_{\text{Exc}}$. See discussion in Appendix B.
@@
Appendix F Relations between the estimates
@@
Let us start with $\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)$. Using results of (Pfau, 2013), we have:
$\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)=\mathbb{E}_{p(\theta\mid D_{tr})}\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\eta_{\tilde{\theta}}\|\eta_{\theta}\Bigr{)}=\\ \mathbb{E}_{p(\theta\mid D_{tr})}D_{G}\Bigl{(}\hat{\eta}_{D_{tr}}\|\eta_{\theta}\Bigr{)}+\mathbb{E}_{p(\tilde{\theta}\mid D_{tr})}D_{G}\Bigl{(}\eta_{\tilde{\theta}}\|\hat{\eta}_{D_{tr}}\Bigr{)}=\tilde{\text{R}}_{\text{Exc}}^{(2,1)}(x)+\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x).$
@@
$\tilde{\text{R}}_{\text{Exc}}^{(1,1)}(x)\geq\tilde{\text{R}}_{\text{Exc}}^{(1,2)}(x)\geq\tilde{\text{R}}_{\text{Exc}}^{(2,2)}(x)=0,$